Consider a rubber ball freely falling from a height h = 4.9 m onto a horizontal elastic plate. Assume that the duration of collision is negligible and the collision with the plate is totally elastic. Then the velocity as a function of time and the height as a function of time will be ?

Grade:12

1 Answers

Saurabh Koranglekar
3 years ago
v = u + gt. As the ball is dropped, v = gt when coming down. v increases makes collision, the value of v becomes +ve, decreases comes to zero and increases. The change from +v to –v is almost instantaneous. Using –ve signs when coming down, (c) is correct. Further h = (1/2)gt2is a parabola.